OpenStudy (anonymous):

What is the product of 6the square root of 7. • 2the square root of 21.? Simplify if possible. Choices: 8the square root of 21. 12the square root of 3. 12the square root of 21. 84the square root of 3.

OpenStudy (anonymous):

from your options you want \[6\sqrt{7}*2\sqrt{3}=12\sqrt{7*3}=?\]

OpenStudy (anonymous):

@undeadknight26 could you help me?

OpenStudy (anonymous):

you can multiply any two real numbers under the radical sign ,provided both are not negative. \[12 \sqrt{7*3}=12 \sqrt{21}\]